OpenStudy (howard-wolowitz):

Can someone help me here:

OpenStudy (howard-wolowitz):

@Michele_Laino can you give me that formula again

OpenStudy (michele_laino):

here is my formula: \[\Large L = \alpha \times R\]

OpenStudy (michele_laino):

you have to transform your angle from degrees to radians, first

OpenStudy (howard-wolowitz):



OpenStudy (michele_laino):

we have to solve this proportion: 300:x=180: pi so we have: x=300*pi/180= 5*pi/3, so it is right. Now you have to apply my formula above \[L = \alpha \times R\]

OpenStudy (howard-wolowitz):

L = ? * ?

OpenStudy (michele_laino):

L=5*(pi/3)*2=...?

OpenStudy (anonymous):

can someone help me?

OpenStudy (michele_laino):

so waht is the right option?

OpenStudy (michele_laino):

@Howard-Wolowitz